OpenStudy (anonymous):

Factor s^3-4s^2-9s+36=

OpenStudy (anonymous):

factor by grouping

OpenStudy (anonymous):

s^2 ( s-4) -9 (s-4) = (s^2 - 9) (s-4) = (s+3)(s-3)(s-4)